OpenStudy (anonymous):

Daniela teaches Ballet. She charges a one−time processing fee of $3.00 and the cost of each class is shown below. Let x represent the number of classes and y represent the cost of classes. Based on this information, what will it cost, in dollars, to take 6 classes? (Do not put a $ in your answer.) Cost of classes (not including processing fee) Number of Classes, x 1 2 3 4 Cost of Classes, y $16.00 $30.00 $56.00 $108.00

OpenStudy (anonymous):

okay, can someone please help explain this to me? Ive become so confused trying to figure it out.

OpenStudy (anonymous):

1. A punter kicks a football upward with an initial speed of 48 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the initial speed. 2. What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)? 3. Given the function g(x) = 8x − 2, find g(−2). I also need help with these three. Please, if anyone can help, please, I really need it!

OpenStudy (anonymous):

you can ignore the three on the bottom for now, I just really need help with the first one

OpenStudy (anonymous):

Are you sure the prices are correct? I found this http://openstudy.com/study#/updates/4f52be31e4b019d0ebb05997 very similar question, but the prices differ. Hmm

OpenStudy (anonymous):

yeah, im sure. I got 143 as my answer for this one though...

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

What did you do to get that answer? :-)

OpenStudy (anonymous):

108+16+16=140 140+3=143?

OpenStudy (anonymous):

thats what the others guy did to get his, but idk if its the correct way to solve this?

OpenStudy (anonymous):

Hmm, but what if we take 3 blocks of 2 classes? That would be 90$ + 3 = 93. For the other problem, the prices were in progression. This does not seem to be the case.

OpenStudy (anonymous):

ohhh, okay, so then how would we solve this one?

OpenStudy (anonymous):

I should probably note that I have a time limit, I hate to rush but its getting close to the end...

OpenStudy (anonymous):

Do you have answer choices?

OpenStudy (anonymous):

no, it just has a blank box to write in...

OpenStudy (anonymous):

I wil keep thinking, but if anything, try your answer.

OpenStudy (anonymous):

if I did, im sure this would be much easier, lol

OpenStudy (anonymous):

okay, 143 is going in the box... lol :)

OpenStudy (anonymous):

could you possibly help with the others now?

OpenStudy (anonymous):

Well, for the punt, you have the height (0) and the initial speed. Solve the quadratic for t :-)

OpenStudy (anonymous):

And pick the largest t value :-)

OpenStudy (anonymous):

okay: -16t^2 + 48t = 0 Factor out -16t -16t(t + 3) = 0 Two solutions t = 0, when it begins it's upward journey and t = 3 sec, when it returns to earth so t = 3 is the answer?

OpenStudy (anonymous):

For this one: Given the function g(x) = 8x − 2, find g(−2). Plug in x = -2 and do the arightmetic.

OpenStudy (anonymous):

Yup, seems correct. Let me check it.

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

and for the second one I got g = 1-4x? is that right?

OpenStudy (anonymous):

No, it should be a number: g(-2) = 8(-2) - 2.

OpenStudy (anonymous):

not the second one, the third one, sorry, lol

OpenStudy (anonymous):

Do you mean the line?

OpenStudy (anonymous):

ohh hold on one sec

OpenStudy (anonymous):

g = 9?

OpenStudy (anonymous):

y=3x-2 is the equation for the line that passes through those two points, now onto finding a parallel that passes through (-2,3) And no, it's -16 -2 = -18 :-)

OpenStudy (anonymous):

The equation is y = 3x + 9.

OpenStudy (anonymous):

so the slope would be 3 right?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

yay! awesome okay now let me just clarify: 1. 3 2. 3 3. g = -18 right

OpenStudy (anonymous):

g(-2) = -18, but I think you only have to enter -18.

OpenStudy (anonymous):

okay, ill just do that, btw, could you also check two problems for me?

OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−3, −1) and (1, 7)? My answer for this is (0,5), is that right? 19. What is the slope of the line passing through the point (4, −3) and perpendicular to the line y = 2x + 1? Write your answer as a fraction. I got (-1/2)x - 1 as my answer for this one, is that correct?

OpenStudy (anonymous):

This one also: Tyrone has a total of 50 DVDs and CD's. If the number of CDs is two more than three times the number of DVDs, how many CDs does he have? I got 14

OpenStudy (anonymous):

ill brb i have to go to the bathroom but please stick around I have some more questions that i want to ask you

OpenStudy (anonymous):

im back, lol

OpenStudy (anonymous):

The first one is correct. For the last one A + B = 50 3A + 2 = B 4A + 2 = 50 A = 48/4 = 12 = Number of DVD's 50 - 12 = 38, the number of CD's

OpenStudy (anonymous):

And I have to go, I am sorry. Have to study for my exam later today. Cya and good luck :-)

OpenStudy (anonymous):

okay, thanks for all of your help

OpenStudy (anonymous):

bye bye and good luck on your exam!